# A multi-centre, multi-device benchmark dataset for landmark-based comprehensive fetal biometry

### v1

@@ Results
We performed a comprehensive evaluation across all three datasets and the combined multi-centre dataset ( MULTI-CENTRE (M-C)). Table 2 reports the Normalised Mean Error (NME) with standard deviation, demonstrating expected domain shift across imaging devices and clinical sites.
@@
Figure 3 shows Bland–Altman plots for each dataset, illustrating the agreement between predicted and ground-truth measurements as a percent difference relative to the ground-truth value. Expressing errors as percentages mitigates heteroscedasticity with increasing measurement size and enables comparison across anatomies. Solid lines indicate the mean bias and dashed lines indicate the 95% limits of agreement (bias 1.96 SD). Distances are computed in image pixel space before conversion to millimetres to avoid confounding effects from dataset-specific calibration.
Figure 4 summarises absolute biometry error (mm) on the M-C test set for models trained on FP, HC18, UCL, and M-C datasets for Head (BPD, OFD), Abdomen (TAD, APAD), and Femur (FL); the y-axis is truncated at 30 mm to improve visualisation of the central distribution. The M-C-trained model achieves competitive medians across all anatomies with consistently narrow interquartile ranges, indicating robust and stable performance. In contrast, HC18-trained models exhibit wider error distributions on the M-C test set, reflecting residual domain shift despite the corrected preprocessing. For head, the median errors of the M-C model (approximately 0.2 mm for BPD and OFD) correspond to clinically acceptable GA variation in late pregnancy.
@@
Annotations are provided as CSV files, each containing landmark coordinates for a single measurement. All CSV files include common fields index (sequential index in the dataset), image_name (filename of the corresponding US image), scale (scaling factor applied during preprocessing, center_w, center_h (centre coordinates of the region of interest, px_to_mm_rate (pixel-to-millimeter conversion rate) mm_dist (distance marker value visible in the image, typically 5mm or 10mm), Algo (device or algorithm identifier), SubjectID (de-identified subject identifier), Split (data split indicator, train or test).
Head Measurements (BPD and OFD): ofd_1_x, ofd_1_y, ofd_2_x, ofd_2_y (landmarks for occipito-frontal diameter) and ofd_1_x, ofd_1_y, ofd_2_x, ofd_2_y (landmarks for bi-parietal diameter)
@@ Discussion
We present the first comprehensive, open-access, multi-centre dataset for landmark-based fetal biometry automation. By combining three sources (FP, HC18, UCL) across 7 US devices and 3 clinical sites, we provide 4,513 annotated images that capture realistic clinical variability. Unlike single-anatomy (HC18) or single-site (AutoFB) datasets, ours enables systematic quantification of domain shift, a critical barrier to clinical deployment of artificial intelligence (AI)-based biometry systems. Our dataset and experiments focus on landmark detection within pre-validated SPs. The variability introduced by SP selection was not quantified in this work. While this approach isolates variability in landmark placement for rigorous benchmarking, a fully automated pipeline would need to address either plane verification or robustness to plane variations. Future work should extend benchmarking to include raw ultrasound sweeps or near-boundary planes to quantify SP selection variability and provide a pathway toward fully autonomous fetal biometry.
@@
Before correcting the preprocessing, the HC18 dataset showed markedly inferior cross-domain generalisation due to landmark-space mismatch. After applying image-centric preprocessing parameters (Section Methods), HC18-trained models show competitive within-domain performance and moderate cross-domain degradation (Table 2: HC18FP = 0.060.07, HC18UCL = 0.150.16). Residual differences in annotation methodology remain due to differing annotation procedures, underscoring the importance of harmonising annotation protocols in multi-centre dataset development.
These findings have important implications for clinical deployment. The marked performance degradation from within-domain to cross-domain settings indicates that models trained on single-site data are unlikely to maintain accuracy when deployed in different clinical settings with varying US devices and operator practices. The improved cross-site generalisation of multi-centre models (e.g., M-CUCL achieving NME values 4–8 lower than FPUCL for head biometry) suggests that robust clinical deployment will require training on diverse, multi-source datasets that capture real-world variability. For sonographers, successful automation could reduce measurement time and inter-operator variability (currently 4.9–11.1% across observers) [sarris2012intra], but only if models are validated across the full spectrum of clinical acquisition conditions encountered in practice.
@@ Methods
| Dataset | Source/Institution | #Subjects (Pregnancies) | #Images | Anatomical Planes | Annotation Tool | Notes |
@@
After acquisition and annotation, all images underwent standardised preprocessing and annotation structure to ensure consistency and reproducibility across different US devices and acquisition protocols.
@@
Manual landmark annotation (FP and UCL): Expert sonographers manually placed anatomical landmarks using the VIA tool [dutta2019vgg], following ISUOG guidelines [salomon2019isuog]. FP annotations were performed by senior maternal-fetal clinicians (inter-rater agreement 93–95% [burgos2020zdataset]); UCL annotations were performed by a clinical research fellow under the supervision of a senior consultant. Each measurement requires two points marking the start and end of the biometric diameter; on average, each plane annotation for landmarks took 20 seconds, which is lower than the 70 seconds required for manual structure segmentation delineation [bano2021autofb]. [bano2021autofb].
@@
For data augmentation, random transformations were applied on the fly during training to increase robustness to operator- and device-related variations. These included random rotations within , horizontal and vertical flips, random scaling (90–110%), translation (up to 10% of the image size), and brightness/contrast jitter (up to 20%). All augmentation parameters and intensity normalisation follow the settings reported in AutoFB [bano2021autofb], which demonstrated effective regularisation for heterogeneous US data.
@@ Acknowledgements
This work was supported in whole, or in part, by the Wellcome/EPSRC Centre for Interventional and Surgical Sciences (WEISS) [203145/Z/16/Z], the Engineering and Physical Sciences Research Council (EPSRC) [EP/P027938/1, EP/R004080/1, EP/P012841/1], the Royal Academy of Engineering under the Chair in Emerging Technologies programme, the Horizon 2020 FET Open [863146], and the Kamin Grants [63418, 72126] from the Israel Innovation Authority. We acknowledge the support of D. Casagrandi and P. Pandya in acquiring the data used in this study.
@@
Accession codes: All data and annotations are available on the UCL Research Data Repository. Taining code and evaluation pipelines are available at https://github.com/surgical-vision/Multicentre-Fetal-Biometry.git.

### v2

@@ Results
We performed a comprehensive evaluation across all three datasets and the combined multi-centre dataset ( MULTI-CENTRE (M-C)). For each measurement, localisation error is quantified using the Normalised Mean Error (NME) in image space: , where and are the sums of endpoint errors under standard and swapped correspondences, respectively, making the metric invariant to endpoint ordering. Table 2 reports the mean and standard deviation of across the test set, demonstrating expected domain shift across imaging devices and clinical sites.
@@
Figure 3 shows normalised Bland–Altman plots for each dataset, illustrating the agreement between predicted and ground-truth measurements as a percent difference relative to the ground-truth. Expressing errors as percentages mitigates heteroscedasticity with increasing measurement size and enables comparison across anatomies. Solid lines indicate the mean bias and dashed lines indicate the 95% limits of agreement (bias 1.96 SD). Distances are converted from pixels to millimetres before computing mean measurements and percentage differences.
Figure 4 shows absolute biometry error (mm) on the M-C test set for models trained on FP, HC18, UCL, and M-C for head, abdomen, and femur (FL); to focus on the central error distribution, the y-axis is truncated at 30 mm; a small number of larger outliers fall above this range. The M-C-trained model achieves competitive medians across all anatomies with consistently narrow interquartile ranges, indicating robust and stable performance. In contrast, HC18-trained models exhibit wider error distributions on the M-C test set, reflecting residual domain shift despite the corrected preprocessing. For head, the median errors of the M-C model ( 0.2 mm for BPD and OFD) correspond to clinically acceptable GA variation in late pregnancy.
@@
Annotations are provided as CSV files, each containing landmark coordinates for a single measurement. All CSV files include common fields index (sequential index in the dataset), image_name (filename of the corresponding US image), scale (scaling factor applied during preprocessing), center_w, center_h (centre coordinates of the region of interest), px_to_mm_rate (pixel-to-millimetre conversion rate), mm_dist (distance marker value visible in the image, typically 5mm or 10mm), Algo (device or algorithm identifier), SubjectID (de-identified subject identifier), Split (data split indicator, train or test).
Head Measurements (BPD and OFD): bpd_1_x, bpd_1_y, bpd_2_x, bpd_2_y (landmarks for bi-parietal diameter) and ofd_1_x, ofd_1_y, ofd_2_x, ofd_2_y (landmarks for occipito-frontal diameter).
@@ Discussion
We present the first comprehensive, open-access, multi-centre dataset for landmark-based fetal biometry automation. By combining three sources (FP, HC18, UCL) across 7 US devices and 3 clinical sites, we provide 4,513 annotated images that capture realistic clinical variability. Unlike single-anatomy (HC18) or single-site (AutoFB) datasets, our dataset enables systematic quantification of domain shift, a critical barrier to clinical deployment of artificial intelligence (AI)-based biometry systems. Our dataset and experiments focus on landmark detection within pre-validated SPs. The variability introduced by SP selection was not quantified in this work. While this approach isolates variability in landmark placement for rigorous benchmarking, a fully automated pipeline would need to address either plane verification or robustness to plane variations. Future work should extend benchmarking to include raw ultrasound sweeps or near-boundary planes to quantify SP selection variability and provide a pathway toward fully autonomous fetal biometry.
@@
Before correcting the preprocessing, the HC18 dataset showed markedly inferior cross-domain generalisation due to landmark-space mismatch. After applying image-centric preprocessing parameters, HC18-trained models show competitive within-domain performance and moderate cross-domain degradation (Table 2: HC18FP = 0.060.07, HC18UCL = 0.150.16). Residual differences in annotation methodology remain due to differing annotation procedures, underscoring the importance of harmonising annotation protocols in multi-centre dataset development.
These findings have important implications for clinical deployment. The marked performance degradation from within-domain to cross-domain settings indicates that models trained on single-site data are unlikely to maintain accuracy when deployed in different clinical settings with varying US devices and operator practices. The improved cross-site generalisation of multi-centre models (e.g., M-CUCL, achieving substantially lower NME values than FPUCL for head biometry) suggests that robust clinical deployment will require training on diverse, multi-source datasets that capture real-world variability. For sonographers, successful automation could reduce measurement time and inter-operator variability (currently 4.9–11.1% across observers) [sarris2012intra], but only if models are validated across the full spectrum of clinical acquisition conditions encountered in practice.
@@ Methods
| Dataset | Source/Institution | #Subjects | #Images | Anatomical Planes | Annotation Tool | Notes |
@@
After acquisition and annotation, all images underwent standardised preprocessing to ensure consistency and reproducibility across different US devices and acquisition protocols.
@@
Manual landmark annotation (FP and UCL): Expert sonographers manually placed anatomical landmarks using the VIA tool [dutta2019vgg], following ISUOG guidelines [salomon2019isuog]. FP annotations were performed by senior maternal-fetal clinicians (inter-rater agreement 93–95% [burgos2020zdataset]); UCL annotations were performed by a clinical research fellow under the supervision of a senior consultant. Each measurement requires two points marking the start and end of the biometric diameter; on average, each plane annotation for landmarks took 20 seconds, which is lower than the 70 seconds required for manual structure segmentation delineation [bano2021autofb].
@@
For data augmentation, random transformations were applied during training to increase robustness to operator- and device-related variations. These included random rotations within , horizontal and vertical flips, random scaling (90–110%), translation (up to 10% of the image size), and brightness/contrast jitter (up to 20%). All augmentation parameters and intensity normalisation follow the settings reported in AutoFB [bano2021autofb], which demonstrated effective regularisation for heterogeneous US data.
@@ Acknowledgements
This work was supported in whole, or in part, by the Wellcome/EPSRC Centre for Interventional and Surgical Sciences [203145/Z/16/Z], the Engineering and Physical Sciences Research Council (EPSRC) [EP/P027938/1, EP/R004080/1, EP/P012841/1, NS/A000027/1], Wellcome [WT101957], the Royal Academy of Engineering under the Chair in Emerging Technologies programme (CiET1819/2/36), the Horizon 2020 FET Open [863146], and the Kamin Grants [63418, 72126] from the Israel Innovation Authority. We acknowledge the support of D. Casagrandi and P. Pandya in acquiring the data used in this study.
@@
Accession codes: All data and annotations are available on the UCL Research Data Repository. Training code and evaluation pipelines are available at https://github.com/surgical-vision/Multicentre-Fetal-Biometry.git.